OpenStudy (anonymous):

Solve this: log (6x+25) base 5 = log (3x+2) base 5 Step by step solutions pls, thank you!

jimthompson5910 (jim_thompson5910):

Hint: if \[\Large \log_b(x)=\log_b(y)\] then \[\Large x = y\]

OpenStudy (anonymous):

\[\log_{5} (6x+25) =\log_{5} (3x+2)\] \[=>(\log(6 x+25))/(\log(5)) = (\log(3 x+2))/(\log(5))\]\[=> \log(6x+25) =\log (3x+2) \] taking exponential on both sides =>6x-3x=-25+2 =>3x=-23 =>x=\[-23/3\]